# ESG Rating Disagreement and Corporate Total Factor Productivity:Inference and Prediction

### v3

@@ Abstract
This paper examines how ESG rating disagreement ($Dis$) affects corporate total factor productivity (TFP) in China based on data of A-share listed companies from 2015 to 2022. We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, and low-pollution firms. Mechanism analysis shows that green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints. Furthermore, XGBoost regression demonstrates that $Dis$ plays a significant role in predicting TFP, with SHAP showing that the dampening effect of ESG rating disagreement on TFP is still pronounced in firms with large $Dis$ values.
@@
Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Machine learning
@@ 2 Theoretical Analysis and Research Hypotheses
According to information asymmetry theory, ESG rating disagreements can confuse stakeholders and contribute to the degree of information asymmetry between companies and investors. ESG ratings, as non-financial information, provide stakeholders with a more comprehensive insight into companies, thereby alleviating information asymmetry(Xiaohong and Zhenghan, 2023). Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022). Thereby, increasing financing constraints and ultimately resulting in a decline in total factor productivity. Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders. For example, companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in total factor productivity(Peng and Xie, 2024).
@@ 3.1 Sample Selection
We utilizes financial data of A-share listed companies in China from 2015 to 2022111Before 2015, the only more well-known local ESG rating agencies in China were Huazheng, alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and MioTech. The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails. All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 13417 samples in baseline regression.
@@ 3.2 Model Specification
To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react, TFP is lagged by one period:
@@ 3.3 Variable Descriptions
The independent variable in this study is ESG Rating Disagreement. Domestic ESG rating agencies have advantages in terms of time and spatial proximity that foreign agencies cannot match, making them more likely to influence the behavior of domestic firms. Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and MioTech. Following Avramov et al. (2022), the ratings were assigned positive values according to their rating levels, with a difference of 1 between levels. We define the rankings as the values normalized to the range [0, 1] for the given rating agency in a given year, and the pairwise standard deviation of these rankings is used as the measure of $Dis$.
@@ 4 Empirical Analysis
Table 2: Descriptive Statistics of Main Variables
Variables Count Mean SD Min 25% 50% 75% Max
$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
@@ 4.1 Baseline Regression
Table 3 reports the results of the baseline regression, with clustering at the individual and time levels. The results indicate that ESG rating disagreement reduces corporate total factor productivity, providing preliminary support for the validity of Hypothesis H1.
@@
4.2 Robustness Tests
Second, since the total factor productivity of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by Lu et al. (2024) as a control variable (see table9).
Third, following Yang et al. (2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable. A two-stage least squares (2SLS) estimation is then conducted to further mitigate potential endogeneity issues (see table10).
Fourth, we denote disagreements under 20% as non-existent and the remainder as present. Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see table11).
@@ 4.3.1 The Role of Green Innovation
Table 4 and figure1 reports the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. Dose this contradicts exiting researches that show firms tend to increase their green innovation efforts when ESG rating disagreement is present(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
@@ 4.4 Heterogeneity Analysis
Table 6 reports the results of the heterogeneity analysis. The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, and low-pollution firms. The above heterogeneity may exist because, first, there is a greater willingness to engage in ESG construction among non-state-owned firms due to the fact that they tend to focus more on ESG performance to attract more investment; second, usually capital-intensive firms are less affected by market forces due to their own low variable costs and high fixed costs; third, in the case of highly polluting firms that are susceptible to a variety of environmental policies and regulations in their financing, such firms are less likely to invest in ESG performance, and therefore the impact of divergent ESG ratings is less significant.
@@
State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution
@@
$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519**
(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484)
Chow Test 498.97*** 248.15*** 475.42***
$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321
Controls YES YES YES YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Double Clustering YES YES YES YES YES YES
$N$ 4416 9001 2342 11075 3900 9517
@@ 5 Further Discussion
We use 80% of the dataset for training and 20% for testing and conducted three rounds of training: the first round included ESG rating disagreement, the second round added ESG rating disagreement from the previous period, and the third round did not include ESG rating disagreement. Table 7 reports the evaluation metrics for the three rounds of training. It can be observed that using $R^{2}$ as the evaluation metric, the model’s predictive performance improved by 4.3% and 6.23% respectively compared to the model without ESG rating disagreement, indicating that ESG rating disagreement aids in the short-term prediction of $TFP_{LP}$.
@@
Metric Training Set 1 Training Set 2 Training Set 3 Test Set 1 Test Set 2 Test Set 3
MSE 0.2609 0.2274 0.2818 0.3492 0.3300 0.3835
RMSE 0.5108 0.4769 0.5309 0.5909 0.5744 00.6192
MAE 0.3995 0.3729 0.4145 0.4589 0.4523 0.4818
MAPE 4.7786% 4.4569% 4.9457% 5.4649% 5.3644% 5.7293%
R2 0.7685 0.7969 0.7500 0.7045 0.7176 0.6755
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in the predictive task, SHAP was used to interpret the XGBoost model. SHAP values, rooted in game theory, quantify each feature’s contribution to the model output by treating features as "players" in a game. The Shapley value $\phi_{i}$ for a feature $i$ is calculated as:
@@
where $S$ is a subset excluding $i$, $f(S)$ is the model output for $S$, and $f(S\cup\{i\})$ is the output after adding $i$ to $S$.
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting total factor productivity. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s total factor productivity, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
@@
6 Conclusion
ESG rating disagreement poses challenges to the enhancement of corporate TFP. This study utilizes financial data from China’s A-share listed companies and ESG rating information from domestic ESG rating agencies between 2015 and 2022 to explore the relationship between ESG rating disagreements and firms’ TFP. The findings robustly indicate that ESG rating disagreements have a negative impact on firms’ TFP. The mechanism analysis reveals that green innovation instead strengthens the dampening effect of ESG rating disagreement on TFP; moreover, ESG rating disagreement strengthens firms’ financing constraints, thereby hindering the improvement of TFP. Heterogeneity analysis further indicates that the negative impact of ESG rating disagreements on TFP is more pronounced in state-owned enterprises, non-capital-intensive firms, and companies with lower levels of pollution. Additionally, an important finding from the analysis using machine learning methods suggests that ESG rating disagreements serve as a significant indicator for short-term predicting firms’ TFP, providing new insights for companies to engage in ESG risk management to enhance TFP.

### v4

@@ Abstract
This paper examines how ESG rating disagreement ($Dis$) affects corporate total factor productivity (TFP) in China based on data of A-share listed companies from 2015 to 2022. We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, low-pollution and high-tech firms, green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints and weakening human capital. Furthermore, XGBoost regression demonstrates that $Dis$ plays a significant role in predicting TFP, with SHAP showing that the dampening effect of ESG rating disagreement on TFP is still pronounced in firms with large $Dis$ values.
@@
Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Human Capital; Machine learning
@@ 2 Theoretical Analysis and Research Hypotheses
According to information asymmetry theory, ESG rating disagreements can confuse stakeholders and contribute to the degree of information asymmetry between companies and investors. ESG ratings, as non-financial information, provide stakeholders with a more comprehensive insight into companies, thereby alleviating information asymmetry(Xiaohong and Zhenghan, 2023). Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022). Thereby, increasing financing constraints and ultimately resulting in a decline in TFP. Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders. Companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in TFP(Peng and Xie, 2024). Meanwhile, the divergence in ESG ratings may also drive companies to focus on improving their ESG recognition, potentially causing an imbalance between the interests of external investors and internal employees. This imbalance could lead to a human capital outflow, ultimately affecting the development of TFP.
@@ 3.1 Sample Selection
We utilize financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave. The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails. All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 3656 firms and 13417 samples in baseline regression.
@@ 3.2 Model Specification
To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react and mitigate potential reverse causation problems, TFP is lagged by one period.:
@@ 3.3 Variable Descriptions
The independent variable in this study is ESG Rating Disagreement. Domestic ESG rating agencies have advantages in terms of time and spatial proximity that foreign agencies cannot match, making them more likely to influence the behavior of domestic firms. Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave. Following Avramov et al. (2022), the ratings were assigned positive values according to their rating levels, with a difference of 1 between levels. We define the rankings as the values normalized to the range [0, 1] for the given rating agency in a given year, and the pairwise standard deviation of these rankings is used as the measure of $Dis$.
@@
$HC$ CSAMR Natural logarithm of number of graduate students and above and natural logarithm of number of undergraduate students and above
@@ 4.1 Baseline Regression
Table 2 reports the results of the baseline regression, with clustering at the individual and time levels. The results indicate that ESG rating disagreement reduces corporate TFP, providing preliminary support for the validity of Hypothesis H1.
@@ 4.2.1 The Role of Green Innovation
Table 3 and figure1 reports the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. The result seem contradict exiting researches, which argue ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
@@
To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the tabel 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
Table 4: Heterogeneity of Green Innovation Mechanism
Variables TFP
High $Dis$ Low $Dis$
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$ 0.0527 0.0616 0.0501 -0.2268*** -0.1532** -0.1554**
(1.2360) (1.5631) (1.2368) (-2.8516) (-2.0471) (-1.9903)
$EI$ 0.0364*** 0.0280*** 0.0230** 0.0092 0.0043 0.0041
(3.1495) (2.6193) (2.0914) (1.4601) (0.7258) (0.6628)
$Dis*EI$0.0522** -0.0461** -0.0381** 0.0433 0.0280 0.0260
(-2.1556) (-2.0558) (-1.6525) (1.2832) (0.8798) (0.7855)
$TobinQ$0.0009 -0.0012 0.0020 -0.0025 -0.0020 -0.0013
(-0.1395) (0.1681) (0.3447) (-0.4475) (-0.3853) (-0.2318)
$Top10$0.0008 -0.0012 -0.0005 -0.0001 0.0004 0.0008
(-0.8389) (-1.3029) (-0.5405) (-0.1141) (0.4905) (0.8934)
$ListAge$ 0.1137** 0.0833* 0.0257 0.0609 0.0529 -0.0332
(2.2567) (1.7886) (0.5359) (1.4289) (1.3186) (-0.7922)
$Lev$0.1140 -0.0997 -0.0976 -0.2786*** -0.2703*** -0.2712***
(0.1716) (-1.2930) (-1.2306) (-3.2945) (-3.3962) (-3.2670)
$Size$ 0.3607*** 0.2525*** 0.2173*** 0.3523*** 0.2355*** 0.2123***
(16.816) (12.732) (10.655) (17.502) (12.434) (10.746)
$Avg$ 0.2126** 0.2424*** 0.2447*** 0.2301*** 0.2706*** 0.2651***
(13.198) (16.277) (15.983) (13.969) (17.451) (16.389)
$Der$ 0.0224** -$0*$0226** 0.0046 0.0052 0.0036
(-2.2377) (-2.2735) (-2.3751) (0.3638) (0.4404) (0.2910)
Controls YES YES YES YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
$N$ 6774 6774 6774 6643 6643 6643
$R^{2}$ 0.1578 0.1434 0.1229 0.1717 0.1533 0.1310
@@ 4.2.2 The Role of Financing Constraints
4.2.3 The Role of Human Capital
Human capital is widely acknowledged as a positive contributor to TFP(Miller and Upadhyay, 2000; Xu et al., 2008; Huang et al., 2024). Employees with a bachelor’s or master’s degree typically reflects higher levels of knowledge, skills, and expertise among workers, enabling them to perform task more effciently, reduce errors, and enhance productivity. In this subsection, we delve deeper into how ESG rating disagreement interacts with human capital and its potential impact on a company’s TFP, with a specific focus on employees holding bachelor’s or master’s degrees. Building on Jiang (2022)’s research on the transmission mechanism, and following a similar approach to verifying the financing constraint mechanism to avoid endogeneity issues in mediating variables, we perform regressions of $Dis$ on $Master$ and $Dis$ on $Undergrad$.
Table 6 reports the regression results, showing that Dis is detrimental to the development of firms’ human capital, which serves as a crucial driver of TFP. Based on this evidence, we argue that Dis reduces TFP by impairing human capital.
Table 6: Human Capital Mechanism
Variables Human Capital
$Master$ $Master$ $Undergrad$ $Undergrad$ $Master$ $Master$ $Undergrad$ $Undergrad$
$Dis$0.0595*** -0.0526*** -0.0345** -0.0316**
(-2.8695) (-3.1495) (-1.7625) (-2.0461)
$Disb$0.0598*** -0.0528** -0.0322* -0.0248*
(-2.8867) (-3.1608) (-1.8268) (-1.9302)
$TobinQ$ 0.0197*** -0.0067 0.2724*** -0.0036
(5.0228) (-1.4155) (19.528) (-0.9603)
$Top10$ 0.0029*** 0.0008 0.0011 0.0003
(4.4620) (1.0079) (0.4783) (0.4613)
$ListAge$0.0010 0.1365*** 0.5530*** 0.1578***
(-0.0413) (3.7156) (6.5333) (5.5267)
$Lev$ 0.0536 -0.0620 6.7446*** 0.0894*
(0.8913) (-0.9460) (32.842) (1.7679)
$Size$ 0.6722*** 0.4376*** -0.4512*** 0.4410***
(49.844) (27.508) (-9.5683) (35.103)
$Avg$0.1597*** -0.0907*** -0.5288*** -0.0739***
(-14.054) (-7.0579) (-13.531) (-7.5015)
$Der$0.0256*** -0.0128 0.0191 -0.0184***
(-3.1008) (-1.5259) (0.6820) (-2.8025)
Controls NO NO NO NO YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
$N$ 10461 10461 12595 12595 10461 10461 12595 12595
$R^{2}$ 0.0011 0.0011 0.0010 0.0011 0.2790 0.1145 0.4112 0.1511
  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and human capital in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and human capital in year t.
@@ 4.3 Heterogeneity Analysis
Table 7 reports the results of the heterogeneity analysis. The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, low-pollution firms and high-tech firms. This heterogeneity may arise for the following reasons: first, non-state-owned firms are generally more willing to engage in ESG initiatives as they tend to focus more on ESG performance to attract investment; second, capital-intensive firms are usually less influenced by market forces due to their low variable costs and high fixed costs, which buffer them from the pressures of ESG-related changes; third, highly polluting firms, which are often subject to stringent environmental policies and regulations in their financing, are less likely to invest in ESG performance, making the impact of divergent ESG ratings less pronounced. And, finally, as previously analyzed, ESG rating disagreements can lead to the loss of highly qualified employees, on whom high-tech firms are particularly dependent. Consequently, the negative impact of disagreements on TFP is more significant for high-tech firms.
@@
State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution High-Tech Non-High-Tech
@@
$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519** -0.0413** -0.0122
(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484) (-2.4165) (-0.4205)
Chow Test 498.97*** 248.15*** 475.42*** 698.54***
$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321 0.1256 0.1315
Controls YES YES YES YES YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
Double Clustering YES YES YES YES YES YES YES YES
$N$ 4416 9001 2342 11075 3900 9517 8454 4963
@@ 5 Further Discussion
We use 80% of the dataset for training and 20% for testing, conducting three rounds of training. The first round includes ESG rating disagreement (Dis), the second round incorporates ESG rating disagreement from the previous period, and the third round excludes ESG rating disagreement altogether. Table 8 presents the evaluation metrics for these three rounds of training.
First, compared to the linear model (Set4), XGBoost demonstrates significantly better performance across all metrics, highlighting the machine learning model’s ability to capture nonlinear relationships. Second, when comparing XGBoost without incorporating Dis (Set3), the model’s performance improves by 4.3% (Set1) and 6.23% (Set2), respectively, using $R^{2}$ as the evaluation criterion. This improvement underscores the predictive value of ESG ratings in modeling TFP.
@@
Metric Training Set 1 Training Set 2 Training Set 3 Training Set 4 Test Set 1 Test Set 2 Test Set 3 Test Set 4
MSE 0.2609 0.2274 0.2818 0.4714 0.3492 0.3300 0.3835 0.4728
RMSE 0.5108 0.4769 0.5309 0.6866 0.5909 0.5744 0.6192 0.6876
MAE 0.3995 0.3729 0.4145 0.5383 0.4589 0.4523 0.4818 0.5415
MAPE 4.7786% 4.4569% 4.9457% 6.4047% 5.4649% 5.3644% 5.7293% 6.4399%
R2 0.7685 0.7969 0.7500 0.5794 0.7045 0.7176 0.6755 0.5955
  • Note: Set1 is the XGBoost performance of $Dis$ in the same period as the TFP, Set2 is the XGBoost performance of $Dis$ in the preceding period, Set3 is the XGBoost performance excluding $Dis$, and Set4 is the model performance of the linear model (no fixed effects) and $Dis$ in the preceding period.
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we applied SHAP to interpret the results of the XGBoost model. SHAP values are grounded in cooperative game theory and provide a way to understand each feature’s contribution to the model output by treating features as "players" in a game. The contribution, or Shapley value $\phi_{i}$, for a particular feature $i$ is computed by averaging its marginal contributions across all possible combinations of features.
Mathematically, the Shapley value $\phi_{i}$ for feature $i$ is defined as:
@@
In this equation:
  • $N$ represents the set of all features, while $S$ is a subset of $N$ that does not include feature $i$.
  • $|S|$ denotes the size of subset $S$.
  • $f(S)$ is the model output based on the subset of features $S$.
  • $f(S\cup\{i\})$ is the output when feature $i$ is added to $S$.
The term $\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}$ in the equation represents a weighting factor that accounts for all possible orders in which features can be added to the model. By calculating $\phi_{i}$ across all possible subsets $S$, SHAP captures the average marginal effect of adding feature $i$ to different subsets of features, thereby quantifying the feature’s contribution to the model’s output.
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP [^3] The importance of each feature is calculated as the percentage of the average of the absolute SHAP values from each to the mean of the absolute SHAP values of all features; in short, a higher percentage means that the feature is more important among all features.. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task [^4] The beeswarm plot shows the effect of different values of feature on TFP, the color of the dots in the plot shifts toward red for larger eigenvalues and toward blue for smaller eigenvalues and points positioned on the horizontal axis with values less than 0 indicate a negative effect on the dependent variable in the model, while points with values greater than 0 signify a positive effect.. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
@@
6 Conclusion and Implication
ESG rating disagreements present significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue. Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreements on TFP. Additionally, these disagreements hinder TFP growth by intensifying financing constraints and weakening human capital. To mitigate these impacts, companies should focus on improving the quality of ESG information disclosure, carefully navigate the trade-offs between green innovation, financing needs, and strive to balance the interests of diverse stakeholders. Regulatory authorities, in turn, should strengthen the scrutiny of green innovation patents to ensure their quality and prevent superficial compliance efforts.
Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreements on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises. These entities should pay closer attention to ESG rating discrepancies, while regulators should provide targeted support to low-pollution and high-tech firms to foster sustainable development and technological innovation. Moreover, ESG rating disagreements serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact. This indicator can be leveraged by listed companies and regulatory agencies as a monitoring and early warning tool to identify potential risks to TFP and implement proactive measures.